Tech Meeting Minutes 20100527
Present: Graeme, Mike, Sam, Andrew, Wahid
|Table of contents|
- LHCb upload problems at Glasgow. Dug has confirmed that the lcg-cr fix for dteam still works, so the NAT problem is "solved" (100 transfers, 98% success; 300 83% success, but likely due to RAL endpoint limitations). Can we try this with an LHCb proxy? ACTION: Ask Vladimir. Have applied for VO membership (Mike). LHCb don't seem to send pilots to sites which are full (for
us full of ATLAS). Maybe chase this later.
- LHCb problems at Edinburgh. Ticket 57042 finally closed. Andy will still migrate s/w area to nfs (pending ACTION). Software should be wiped and reinstalled from scratch.
- ATLAS production has restarted slowly at ECDF. ACTION: Andy to keep an eye on this.
- Sam fixed an ILC authentication problem on the Durham SE, so cross-site login still works.
- squid: ECDF fails SAM tests through the Lancaster squid, so there is still a firewall issue. ACTION: Wahid to fix.
- CREAM: Still waiting for extra servers from systems team. Decomissioning old UI is slower than expected.
- Storm + GPFS: ACL problem fixed (no attribute caching mount option). Suggested now to move the ECDF SCRATCHDISK to storm to avoid test endpoints in ToA. To do this:
- Set ANALY_ECDF offline (it's online right now).
- Move datasets off current SCRATCHDISK to Glasgow SCRATCHDISK.
- Delete ECDF SCRATCHDISK copies.
- Change ToA.
- Move user datasets back.
- Set site 'brokeroff' to hammercloud it.
- Set site 'online'.
- Graeme sent email to Jose to test this.
- Glasgow is in cooling off period.
- Graeme has a dead DSL router at home. He's trying to get to Ferney to get France Telecom to replace it. Home internet down for now!
- Dug is leaving at the end of this month, so big thanks to him for all his work in the last 18 months. Thanks!
- Meeting next week: http://indico.cern.ch/conferenceDisplay.py?confId=96377.